# Towards Explainable Spoofed Speech Attribution and Detection: a Probabilistic Approach for Characterizing Speech Synthesizer Components

### v2

@@ 3.4 Explanability with Shapley values
where $S$ represents a subset of features excluding $\rho_{\ell m}$, and $|S|$ denotes the cardinality of $S$. In practice, for any given classifier, the Shapley values are approximated by considering various combinations of the features, including and excluding ($\rho_{\ell m}$, to estimate the contribution of the feature to a specific prediction. The difference between these predictions, representing the marginal contribution of $\rho_{\ell m}$, is then calculated. This process is repeated over numerous such combinations, and the Shapley value for $\rho_{\ell m}$ is obtained by averaging these marginal contributions. In this study, we calculate the Shapley values of attributes $\rho_{\ell m}$ to understand their contributions in performing the downstream tasks. This analysis helps us to identify the key decision-makers in the classification process and quantify their individual contributions to the downstream tasks.
@@ 4.1 Datasets
Table 3: ASVspoof2019-det, ASVspoof2019-attr-2 and ASVspoof2019-attr-17, utterances per attack statistics. F and M correspond to females and males, respectively. #Spk and #Utter: correspond to the number of speakers and utterances, respectively. Blank space signifies no utterance. The same color between ASVspoof2019 and ASVspoof2019 attack attribution signifies the use of the same set of utterances. The eval set of ASVspoof2019-attr-2 protocol used for spoofing attack attribution task only consists of known attacks A16 and A19.
ASVspoof2019-det and ASVspoof2019-attr-2 ASVspoof2019-attr-17
Train Dev Eval Train Dev Eval
#Utter #Spk #Utter #Spk #Utter #Spk #Utter #Spk #Utter #Spk #Utter #Spk
Bonafide 2580 12F,8M 2548 12M,8F 7355 37F,30M
A01 3800 3716 3040 760 3716
A02 3800 3716 3040 760 3716
A03 3800 3716 3040 760 3716
A04 3800 3716 3040 760 3716
A05 3800 3716 3040 760 3716
A06 3800 12F,8M 3716 6F,4M 3040 12F,8M 760 12F,8M 3716 6F,4M
A07 4914 2357 589 1968
A08 4914 2357 589 1968
A09 4914 2357 589 1968
A10 4914 2357 589 1968
A11 4914 2356 590 1968
A12 4914 2357 589 1968
A13 4914 2356 590 1968
A14 4914 2357 589 1968
A15 4914 2357 589 1968
A16 (A04) 4914 2357 589 1968
A17 4914 2357 589 1968
A18 4914 2357 589 1968
A19 (A06) 4914 27F,21M 2356 21F,18M 590 21F,18M 1968 Speaker Common (21F,18M), Speaker Disjoint (6F,3M)
Total (#Spoof) 22800 22296 63882 48878 12220 47880
Total (#Spoof) 108978 108978
@@ 4.2 CM embedding extraction
Considering model generalization, we employ two variations of the AASIST model, RawBoost [75] and an SSL-based model [62], applied on top of the vanilla AASIST model. Firstly, RawBoost-AASIST includes three different data augmentation techniques: (i) linear and non-linear convolutive noise (ii) impulsive signal-dependent additive noise, and (iii) stationary signal-independent additive noise. Linear and nonlinear convolutive noise is introduced to the signal to simulate the effects of encoding, decoding, and compression-decompression processes. Impulsive signal-dependent additive noise mimics the noise caused by improper microphone placement and quantization errors, while stationary signal-independent additive noise simulates the noise introduced by loose or poorly connected cables. All three data augmentation techniques are applied concurrently in our approach. The second model variant, SSL-AASIST, uses both a pre-trained wav2vec 2.0 model [78] (trained in a self-supervised manner) and data augmentation for improved generalization to different domains. It replaces the sinc-layer front-end components with the wav2vec 2.0 model. The model is fine-tuned using in-domain ASVspoof2019 data as pre-training is only conducted over bonafide data.
@@ 4.3 Probabilistic Attribute Extraction
Let us now detail the attribute-value pairs used in developing the attribute extractors. All ASVspoof 2019 utterances were generated using a modular approach. We use the respective metadata available in [63, Table 1] to design and train seven extractors following the attribute-value pairs, as summarized in Table 4. The spoofed utterances including both VC and TTS are generated based on predefined high-level design choices, which include input, input processor, duration model, conversion function, speaker representation, conversion model outputs, and waveform generator (voice coder or vocoder). Note that both ASVspoof2019-det and ASVspoof2019-attr-2 protocols have identical spoofing attacks in the training and development partitions—hence, the derived attribute and value pairs for both protocols are identical. The ASVspoof2019-attr-17 protocol, in turn, contains all the $17$ attacks (note that attack pairs (A04, A16) and (A06, A19) present identical methods). Compared with the two other evaluation protocols, the attribute definitions are identical, but their value sets are now expanded due to the inclusion of more attacks. To sum up, the attribute values are dictated by the metadata available. Note also that the dimensionality of the probabilistic embeddings in ASVspoof2019-attr-17 is twice of that in the two other protocols.
@@ 4.6 Explanability with Shapley values
To gain deeper insights into the decision-making processes of our proposed approach, we conducted a Shapley value [31] analysis to quantify the contribution of each feature (i.e., each attribute value) to the predictions. This analysis was performed using the best-performing back-end classifier for each downstream task. Shapley values were calculated on a per-sample basis, meaning that the contribution of each feature was specific to a particular speech utterance in the evaluation set. The overall contribution of attributes for a given downstream task was determined by averaging the attribute-specific Shapley values across the evaluation set. The ASVspoof2019-det protocol was used for spoofing detection, while the ASVspoof2019-attr-17 protocol was employed for spoofing attack attribution.
@@ 5.1 Performance of original (raw) CM embeddings
While SSL-AASIST shows better performance in spoofing detection, its attribution performance, with a balanced accuracy of $94.09\%$, is lower compared to AASIST. To investigate this, we display attack-specific t-distributed stochastic neighbor embedding (t-SNE) [81] projections of AASIST and SSL-AASIST embedding in Figure 3 on the development data. We make two observations. First, the bonafide samples form distinct clusters in both embedding spaces, demonstrating clear separation from most attacks. Second, the attacks are more overlapped with each other in the SSL-AASIST embedding space. This might be due to the SSL model being pre-trained on a different, out-of-domain dataset. Conversely, the vanilla AASIST is trained solely on ASVspoof2019 and may exhibit less domain mismatch. Based on these findings, we evaluate attack attribution on the larger ASVspoof2019-attr-17 protocol only with the vanilla AASIST embeddings (the last row of Table 6). The highest balanced accuracy of $90.16\%$ and lowest EER of $2.11\%$ are obtained with SVM and logistic regression back-ends. As expected, the overall performance drops on this larger protocol.
@@ 5.2 Performance of Probabilistic Attribute Extractors
With the baseline embedding extractors now being defined, we now turn our attention to the the probabilistic attribute extractors. Despite being discriminatively trained, the attribute extractors are viewed as feature extractors—i.e., part of our model selection strategy. Therefore, all the numbers reported in this subsection correspond to the development parts of the included protocols. The results for the detection of attribute values are shown in Figure 4. Concerning the choice of the CM embeddings, the findings are in line with the attribution performance in Table 6 —namely, SSL-AASIST being inferior (here, by an order of magnitude) to the two other types of embeddings. As for the vanilla vs RawBoost variants of AASIST, the latter is generally inferior (except for the duration and speaker attributes).
As for the relative performance for different attributes using vanilla AASIST, waveform generation achieves the lowest EER of $0.2\%$, followed by input processor ($0.4\%$), outputs ($0.5\%$), inputs ($0.6\%$), conversion ($0.7\%$), duration ($0.9\%$), and speaker ($1.0\%$). In absolute terms all EERs are low ($\leq 1\%$), indicating the potential of a simple ’off-the-shelf’ CM embeddings extractor followed up by MLP in extracting high-level spoofing attributes. Since this model structure is identical across all the attributes, the relative ordering reflects the extent to which traces of these attributes are present in the AASIST CM embeddings. In particular, waveform generator (vocoder) traces are most prominently present.
Figure 4: Probabilistic attribute extractor performance in development set using ASVspoof2019-det protocol for each attribute (ATR) shown in terms of EER.
@@
Figure 5: The Sankey plot of attribute 2 (input processor) with AASIST embeddings in attribute value detection (a) training, (b) development, and (c) evaluation set of ASVspoof2019 attack attribution dataset, LP/MF: LPCC/MFCC, CNN-bRNN: CNN+bi-RNN and MF/i-vec: MFCC/ i-vector. The utterances specific to an attack are assigned to an attribute value based on the maximum prediction score from the probabilistic attribute extractor $\mathcal{F}^{\text{pae}}_{{\bm{\rho}}_{l}}$. This plot can be viewed as an alternative visual representation of the confusion matrix. The visualization provides an interpretation of how well the utterances belonging to a specific attack type can map to their corresponding attribute value using the trained probabilistic attribute extractor.
@@ 5.3 Downstream Tasks using Probabilistic Attribute Embeddings
Overall, how do our proposed probabilistic embeddings compare with the raw CM embeddings? On ASVspoof2019-attr-2, regardless of the auxiliary classifiers, the probabilistic attribute embeddings perform comparably to, or in some cases better than, the raw CM embeddings. We observe similar trend also on the ASVspoof2019-attr-17 protocol (with the DT classifier being an exception). These results confirm that the proposed probabilistic attribute embeddings perform as well as or better than raw CM embeddings, with the added benefits of explainability. It is also noteworthy that the dimensionality of the embeddings used for the downstream tasks is greatly reduced from the original ($160$) to $25$ (in the ‘det’ and ‘attr-2’ protocols) and $50$ (in the ‘attr-17’ protocol). In this sense, our proposed probabilistic achieves better compaction of information between the audio waveforms and the class labels.
@@ 5.4 Explanability of Probabilistic attribute embeddings
As our final analysis, we assess the relative importance of attribute values in probabilistic attribute embeddings to decision-making in downstream tasks using Shapley values [31]. To analyze the relative importance of each attribute value, we selected the best-performing classifier for each downstream task: naive Bayes for the detection and logistic regression for the attribution task. The contributions of these attribute values on the eval-sets are illustrated in Figures 7 (detection) and 7 (attribution). Figure 7 highlights that duration.FF (where “duration” is the attribute and “FF” is a specific value of the attribute), conversion.VAE, and conversion.linear are the top three attribute values predominantly contributing to predictions in the detection task. Similarly, Figure 7 indicates that waveform gen.WORLD, speaker.one-hot, and output.MCC-F0 are the top three attribute values influencing the predictions in the attribution task.
Figure 6: Mean absolute Shapley values for spoofing detection (eval-set of ASVspoof2019-det) using maximum likelihood classifier.
Figure 7: Mean absolute Shapley values for spoofing attack attribution (eval-set of ASVspoof2019-attr-17) using logistic regression classifier.
We observe that, in spoofing detection, duration, conversion, and speaker attributes emerge as key factors in decision-making, suggesting that variations in these attribute values exhibit distinct distributions between spoofed and bonafide utterances, making them highly discriminative. This further indicates that the synthesis process might introduce artifacts linked to these attributes’ design choices, which help differentiate spoofed speech from bonafide speech. Similarly, for spoofing attack attribution, waveform generation, speaker, and output attributes play a crucial role in distinguishing between different spoofing attacks. This suggests that the specific design choices of these attributes might introduce distinct artifacts in the synthesized speech signals, providing valuable cues for effective spoofing attack attribution. These insights can be leveraged in the future to deepen our understanding of the spoofed speech generation process, enabling more transparent and informed decision-making.
@@ 6 A Broader Outlook — Towards Generalizable Spoofed Speech Attributes
The two tasks considered in the present study, spoofing attack detection and attribution, have also been primarily addressed as utterance-level tasks (for exceptions where parts of utterances can be spoofed, see e.g. [93, 51]). This was also our purposefully selected focus, following the mainstream research in deepfake detection and biometric antispoofing. Nonetheless, a spoofed utterance is a time-domain signal where different parts of the utterance could contain different set of artifacts, some of them more informative of specific attributes. Therefore, one clear future direction is to study probabilistic attribution at the level of individual speech segments to obtain a more fine-grained representation, analogous to phonetic posteriorgrams [85] or manner/place extractors [84], but with relevant spoofed speech attributes used instead of phones or distinctive features. As argued in Section 2, however, such explainable characterization of artificial (spoofed) speech does not yet have similarly standardized, de facto attributes. In the present work, we defined the attributes through specific metadata target labels available in a dataset, considered as a ’placeholder’ in the absence of established theory for spoofed speech artifacts. The results presented in the previous section are promising and serve as an initial proof-of-concept. As speech synthesis methodology evolves, it is foreseeable that explainable spoof attributes will likely be tied to specific datasets. The longer-term aim, nonetheless, is to come up with generalizable attributes applicable to a wider array of spoofed speech generators.

### v3

@@ 3.4 Explanability with Shapley values
where $S$ represents a subset of features excluding $\rho_{\ell m}$, and $|S|$ denotes the cardinality of $S$. The function $\mathcal{F}^{BC}$(.) denotes the prediction score output by the classifier. In practice, for any given classifier, the Shapley values are approximated by considering various combinations of the features, including and excluding ($\rho_{\ell m}$), to estimate the contribution of the feature to a specific prediction. The difference between these predictions, representing the marginal contribution of $\rho_{\ell m}$, is then calculated. This process is repeated over numerous such combinations, and the Shapley value for $\rho_{\ell m}$ is obtained by averaging these marginal contributions. In this study, we calculate the Shapley values of attributes $\rho_{\ell m}$ to understand their contributions in performing the downstream tasks. This analysis helps us to identify the key decision-makers in the classification process and quantify their individual contributions to the downstream tasks.
@@ 4.1 Datasets
Table 3: ASVspoof2019-det, ASVspoof2019-attr-2 and ASVspoof2019-attr-17, utterances per attack statistics. F and M correspond to females and males, respectively. #Spk and #Utter: correspond to the number of speakers and utterances, respectively. Blank space signifies no utterance. The same color between ASVspoof2019 and ASVspoof2019 attack attribution signifies the use of the same set of utterances. The evaluation set of the ASVspoof2019-attr-2 protocol used for the spoofing attack attribution task consists only of known attacks A16 and A19. Furthermore, both the ASVspoof2019-attr-2 and ASVspoof2019-attr-17 protocols contain only spoofed utterances and do not include any bonafide speech. In contrast, bonafide utterances are present only in the ASVspoof2019-det protocol, designed for spoofing detection (i.e. bonafide/spoof) classification
. ASVspoof2019-det and ASVspoof2019-attr-2 ASVspoof2019-attr-17 Train Dev Eval Train Dev Eval #Utter #Spk #Utter #Spk #Utter #Spk #Utter #Spk #Utter #Spk #Utter #Spk Bonafide 2580 12F,8M 2548 12M,8F 7355 37F,30M A01 3800 3716 3040 760 3716 A02 3800 3716 3040 760 3716 A03 3800 3716 3040 760 3716 A04 3800 3716 3040 760 3716 A05 3800 3716 3040 760 3716 A06 3800 12F,8M 3716 6F,4M 3040 12F,8M 760 12F,8M 3716 6F,4M A07 4914 2357 589 1968 A08 4914 2357 589 1968 A09 4914 2357 589 1968 A10 4914 2357 589 1968 A11 4914 2356 590 1968 A12 4914 2357 589 1968 A13 4914 2356 590 1968 A14 4914 2357 589 1968 A15 4914 2357 589 1968 A16 (A04) 4914 2357 589 1968 A17 4914 2357 589 1968 A18 4914 2357 589 1968 A19 (A06) 4914 27F,21M 2356 21F,18M 590 21F,18M 1968 Speaker Common (21F,18M), Speaker Disjoint (6F,3M) Total (#Spoof) 22800 22296 63882 48878 12220 47880 Total (#Spoof) 108978 108978
@@ 4.2 CM embedding extraction
Considering model generalization, we employ two variations of the AASIST model, RawBoost [75] and an SSL-based model [62], applied on top of the vanilla AASIST model. Although AASIST-RawBoost is typically associated with improving generalization of spoofing detection to degraded conditions, we include it along with vanilla and SSL variation of AASIST for contrastive purposes, given they have received less attention in ’clean’ conditions and in attribution tasks.
The RawBoost-AASIST includes three different data augmentation techniques: (i) linear and non-linear convolutive noise, (ii) impulsive signal-dependent additive noise, and (iii) stationary signal-independent additive noise. Linear and nonlinear convolutive noise is introduced to the signal to simulate the effects of encoding, decoding, and compression-decompression processes. Impulsive signal-dependent additive noise mimics the noise caused by improper microphone placement and quantization errors, while stationary signal-independent additive noise simulates the noise introduced by loose/poorly connected cables, air conditioning systems, or other electronic sources. All three data augmentation techniques are applied concurrently in our approach. The second model variant, SSL-AASIST, uses both a pre-trained wav2vec 2.0 model [78] (trained in a self-supervised manner) and data augmentation for improved generalization to different domains. It replaces the sinc-layer front-end components with the wav2vec 2.0 model. The model is fine-tuned using in-domain ASVspoof2019 data as pre-training is only conducted over bonafide data.
@@ 4.3 Probabilistic Attribute Extraction
Let us now detail the attribute-value pairs used in developing the attribute extractors. All ASVspoof 2019 utterances were generated using a modular approach. We use the respective metadata available in [63, Table 1] to design and train seven extractors following the attribute-value pairs, as summarized in Table 4. The spoofed utterances including both VC and TTS are generated based on predefined high-level design choices, which include input, input processor, duration model, conversion function, speaker representation, conversion model outputs, and waveform generator (voice coder or vocoder). Note that attribute-value pair characterization depends solely on spoofed speech generation and is independent of bonafide speech. Additionally, both ASVspoof2019-det and ASVspoof2019-attr-2 protocols have identical spoofing attacks in the training and development partitions—hence, the derived attribute and value pairs for both protocols are identical. The ASVspoof2019-attr-17 protocol, in turn, contains all the $17$ attacks (note that attack pairs (A04, A16) and (A06, A19) present identical methods). Compared with the two other evaluation protocols, the attribute definitions are identical, but their value sets are now expanded due to the inclusion of more attacks. To sum up, the attribute values are dictated by the metadata available. Note also that the dimensionality of the probabilistic embeddings in ASVspoof2019-attr-17 is twice of that in the two other protocols.
@@ 4.6 Explanability with Shapley values
To gain deeper insights into the decision-making processes of our proposed approach, we conducted a Shapley value [31] analysis to quantify the contribution of each feature (i.e., each attribute value) to the predictions. This analysis was performed using the best-performing back-end classifier for each downstream task. Shapley values were calculated on a per-sample basis, meaning that the contribution of each feature was specific to a particular speech utterance in the evaluation set. Since Shapley values can be positive (indicating a positive influence on the decision) or negative (indicating a negative influence), both directions contribute meaningfully to the final prediction. Therefore, we consider the absolute Shapley values in our study to measure overall influence, regardless of direction. The contribution of attribute values was determined by ranking the corresponding Shapley values at the utterance level, followed by averaging these rankings across the evaluation set and classes. The ASVspoof2019-det protocol was used for spoofing detection, while the ASVspoof2019-attr-17 protocol was employed for spoofing attack attribution.
@@ 5.1 Performance of original (raw) CM embeddings
While SSL-AASIST shows better performance in spoofing detection, its attribution performance, with a balanced accuracy of $94.09\%$, is lower compared to AASIST. To investigate this, we display attack-specific t-distributed stochastic neighbor embedding (t-SNE) [81] projections of AASIST and SSL-AASIST embedding in Figure 3 on the development data. We make two observations. First, the bonafide samples form distinct clusters in both embedding spaces, demonstrating clear separation from most attacks. Second, the attacks are more overlapped with each other in the SSL-AASIST embedding space. This is likely due to the SSL model being pre-trained on a different, out-of-domain dataset and then fine-tuned on ASVspoof2019 for spoofing detection. While this setup enhances generalization in spoofing detection across diverse attack types, it does dilute attack-specific characteristics. In contrast, AASIST is trained entirely on ASVspoof2019, which helps preserve detailed attack-specific information in its embeddings. Based on these findings, we evaluate attack attribution on the larger ASVspoof2019-attr-17 protocol only with the vanilla AASIST embeddings (the last row of Table 6). The highest balanced accuracy of $90.16\%$ and lowest EER of $2.11\%$ are obtained with SVM and logistic regression back-ends. As expected, the overall performance drops on this larger protocol.
@@ 5.2 Performance of Probabilistic Attribute Extractors
With the baseline embedding extractors now defined, we now turn our attention to the the probabilistic attribute extractors. Despite being discriminatively trained, the attribute extractors are viewed as feature extractors—i.e., part of our model selection strategy. Therefore, all the numbers reported in this subsection correspond to the development parts of the included protocols. The results for the detection of attribute values are shown in Figure 4. Concerning the choice of the CM embeddings, the findings are in line with the attribution performance in Table 6 —namely, SSL-AASIST being inferior (here, by an order of magnitude) to the two other types of embeddings. As for the vanilla vs RawBoost variants of AASIST, the latter is generally inferior (except for the duration and speaker attributes).
As for the relative performance for different attributes using vanilla AASIST, waveform generation achieves the lowest EER of $0.2\%$, followed by input processor ($0.4\%$), outputs ($0.5\%$), inputs ($0.6\%$), conversion ($0.7\%$), duration ($0.9\%$), and speaker ($1.0\%$). In absolute terms all EERs are low ($\leq 1\%$), indicating the potential of a simple ’off-the-shelf’ CM embeddings extractor followed up by MLP in extracting high-level spoofing attributes. Since this model structure is identical across all the attributes, the relative ordering reflects the extent to which traces of these attributes are present in the AASIST CM embeddings. In particular, waveform generator traces are most prominently present.
Figure 4: Probabilistic attribute extractor performance in development set using ASVspoof2019-det protocol for each attribute (ATR) shown in terms of EER, input pro.: input processor and wav. gen.: waveform generator.
@@
Figure 5: The Sankey plot of attribute 2 (input processor) with AASIST embeddings in attribute value detection (a) training, (b) development, and (c) evaluation set of ASVspoof2019 attack attribution dataset, LP/MF: LPCC/MFCC, CNN-bRNN: CNN+bi-RNN and MF/i-vec: MFCC/ i-vector. The utterances specific to an attack are assigned to an attribute value based on the maximum prediction score from the probabilistic attribute extractor $\mathcal{F}^{\text{pae}}_{{\bm{\rho}}_{l}}$. This plot can be viewed as an alternative visual representation of the confusion matrix. The visualization provides an interpretation of how well the utterances belonging to a specific attack type can map to their corresponding attribute value using the trained probabilistic attribute extractor. The corresponding EER for the input processor attribute is (a) $2.7\%$ on the training set, (b) $3.2\%$ on the development set, and (c) $3.1\%$ on the evaluation set.
@@ 5.3 Downstream Tasks using Probabilistic Attribute Embeddings
Overall, how do our proposed probabilistic embeddings compare with the raw CM embeddings? On ASVspoof2019-attr-2, regardless of the auxiliary classifiers, the probabilistic attribute embeddings perform comparably to, or in some cases better than, the raw CM embeddings. We observe similar trend also on the ASVspoof2019-attr-17 protocol (with the DT classifier being an exception). These results confirm that the proposed probabilistic attribute embeddings perform as well as or better than raw CM embeddings, with the added benefits of explainability. It is also noteworthy that the dimensionality of the embeddings used for the downstream tasks is greatly reduced from the original ($160$) to $25$ (in the ‘det’ and ‘attr-2’ protocols) and $50$ (in the ‘attr-17’ protocol). In this sense, our proposed probabilistic approach achieves better compaction of information between the audio waveforms and the class labels.
@@ 5.4 Explanability of Probabilistic attribute embeddings
As our final analysis, we assess the relative importance of attribute values in probabilistic attribute embeddings to decision-making in downstream tasks using Shapley values [31]. To analyze the relative importance of each attribute value, we selected the best-performing classifier for each downstream task: naive Bayes for the detection and logistic regression for the attribution task. The contributions in terms of ranking of these attribute values and attributes on the eval-sets are illustrated in Figure 6 (a, b) for detection and (c, d) for attribution. The overall ranking of each attribute is computed as the average of its individual attribute value rankings.
Figure 6: Average rankings of Shapley value across utterances and classes for: (a) spoofing detection based on the top $10$ attribute values, (b) spoofing detection based on attributes, (c) spoofing attack attribution based on the top $10$ attribute values, and (d) spoofing attack attribution based on attributes. A lower rank indicates a higher contribution to decision-making.
Figure 6 (a) highlights that waveform gen.concat (where ”waveform gen” is the attribute and ”concat” is a specific value), outputs.MCC-F0, and waveform gen.WaveNet are the top three attribute values that predominantly contribute to predictions in the spoofing detection task. When aggregated at the attribute level, waveform gen, outputs, and inputs emerge as the most influential attributes in spoofing detection. Similarly, Figure 6 (c) shows that input.text, input.speech (human), waveform gen.WORLD, and speaker.one-hot are among the top attribute values influencing predictions in the spoofing attack attribution task. In terms of aggregated contribution, the attributes inputs, speaker, and duration are the most dominant in attribution.
We observe that, in spoofing detection, waveform gen, outputs, and inputs attributes emerge as key factors in decision-making, suggesting that variations in these attribute values exhibit distinct distributions between spoofed and bonafide utterances, making them highly discriminative. This further indicates that the synthesis process might introduce artifacts linked to these attributes’ design choices, which help differentiate spoofed speech from bonafide speech. Similarly, for spoofing attack attribution, inputs, speaker, and duration attributes play a crucial role in distinguishing between different spoofing attacks. This suggests that the specific design choices of these attributes might introduce distinct artifacts in the synthesized speech signals, providing valuable cues for effective spoofing attack attribution. These insights can be leveraged in the future to deepen our understanding of the spoofed speech generation process, enabling more transparent and informed decision-making. For further insight, class-wise rankings of Shapley values for all attributes and their values are available on our public GitHub [^11] https://github.com/Manasi2001/Spoofed-Speech-Attribution/tree/main/SHAP-plots repository. We warmly encourage interested readers to explore these resources for a more detailed understanding.
@@ 6 A Broader Outlook — Towards Generalizable Spoofed Speech Attributes
The two tasks considered in the present study, spoofing attack detection and attribution, have also been primarily addressed as utterance-level tasks (for exceptions where parts of utterances can be spoofed, see e.g. [93, 51]). This was also our purposefully selected focus, following the mainstream research in deepfake detection and biometric antispoofing. Nonetheless, a spoofed utterance is a time-domain signal where different parts of the utterance could contain different set of artifacts, some of them more informative of specific attributes. Therefore, one clear future direction is to study probabilistic attribution at the level of individual speech segments to obtain a more fine-grained representation, analogous to phonetic posteriorgrams [85] or manner/place extractors [84], but with relevant spoofed speech attributes used instead of phones or distinctive features. As argued in Section 2, however, such explainable characterization of artificial (spoofed) speech does not yet have similarly standardized, de facto attributes. In the present work, we defined the attributes through specific metadata target labels available in a dataset, considered as a ’placeholder’ in the absence of established theory for spoofed speech artifacts. The results presented in the previous section are promising and serve as an initial proof-of-concept. Readers interested in understanding how the probabilistic attributes behave when presented with bonafide speech samples are referred to the Appendix-2. As speech synthesis methodology evolves, it is foreseeable that explainable spoof attributes will likely be tied to specific datasets. The impact of unknown spoofing attacks on the proposed explainable framework is discussed in Appendix-3. The longer-term aim, nonetheless, is to come up with generalizable attributes applicable to a wider array of spoofed speech generators.